OpenStudy (anonymous):

In trapezoid ABCD, if the base DC = 16 and the base AB = 20 and E is the midpoint of AD and F is the midpoint of BC, what is the measure of EF? 1) 36 2) 18 3) 52 4) not enough information

OpenStudy (anonymous):

18

OpenStudy (anonymous):

this = (AB + CD) / 2 = (20+16) /2 = 18